What Transformation of y=2^x results in the equation 1/5(y-3)=2^-(x-4)? Is it:
a)horizontal translation of 4,vertical translation of 3,stretch of 5, reflection in x-axis
b) ht of -4,vt of 3, stretch of 5, reflection in x-axis
c) ht of 4,vt of 3, stretch of 5, reflection in y-axis
d) ht of -4,vt of 3, stretch of 5, reflection in y-axis
Can someone please help me??????

Answer by stanbon(75887)

What Transformation of y=2^x results in the equation 1/5(y-3)=2^-(x-4)? Is it:
a)horizontal translation of 4,vertical translation of 3,stretch of 5, reflection in x-axis
b) ht of -4,vt of 3, stretch of 5, reflection in x-axis
c) ht of 4,vt of 3, stretch of 5, reflection in y-axis
d) ht of -4,vt of 3, stretch of 5, reflection in y-axis
Can someone please help me??????
-----------------------
What Transformation of y=2^x results in the equation 1/5(y-3)=2^-(x-4)?
Rearranging:
y = 5(2^(-(x-4))+3
--------
You can see there is a stretch of 5, there is a reflection in
he y-axis because of the "-"in front of (x-4).
But I don't know what you mean by "ht".
I do know there is a traslation of 4 to the right.
Does that mean ht is 4 or ht is -4?
--------------
If 4 to the right means ht=4, c is the answer
If 4 to the right mean ht=-4, d is the answer.
===============
Maybe you would oblige me and let me know exactly what ht=4 means.
================
Cheers,
Stan H.
Cheers,
Stan H.